OpenStudy (anonymous):

if f(x) = 3x+2 and g(x) = x^2 +4, what is f(g(x))?

OpenStudy (anonymous):

3(x^2+4)+2 <--- you get f(x) and just plug in the function of g(x) for every x you see in the f(x) equation 3x^2+12+2 <--- distribute the 3 3x^2+14 <--- add the whole numbers 3x^2+14=0 <--- set it equal to zero and get as an answer ... x=squareroot -i(14/3) or x=squareroot i(14/3) <--- cant have a negative under the squareroot that's why we have an imaginary number Hope this helps :)

jhonyy9 (jhonyy9):

@mevargas ATTENTION ! 3x^2 +14 =0 3x^2 = -14 x^2 = - 14/3 x_1,2 = +/- isqrt(14/3) = so what you can continue again ratzionalizing the denominator